# Towards AI Accountability Infrastructure: Gaps and Opportunities in AI Audit Tooling

### v1

@@ 5.2. Moving beyond ad hoc toolkits, towards common infrastructure
Our determination of the performance of the algorithm carries a lot of weight within the organization. It’s not like somebody else could just throw it and be like “Oh, we’ll go ask somebody else”… because we built the infrastructure [to test algorithms], so we have that lever. So I almost feel like, if you can control the data sources or the ways to integrate algorithms into the data sources, that gives you power. (P3)
Toolkits currently dominating the audit tooling landscape are devoted mostly to performance evaluation and were built mostly by large, for-profit tech companies (Figures 2 & C.7). Especially the external auditors interviewed seemed to be either skeptical of the data provided from these tools or aware of their unreliability, naming the shutdowns of Reddit and Twitter’s APIs (P24) as well as Facebook’s CrowdTangle tool (P12) (see Section  4.3). Open-source tools can often fill the gap, however, due to the frequent context-specific customizations required for such a varied and multi-disciplinary user base, many tools end up suitable only for a constrained set of methods or targets. As one government official noted, “Some methods might be appropriate for real-time systems. Some methods might be appropriate for after-the-case assessment, as some might need case-by-case analysis, and some might be applicable to the entire training or assessment data” (P21). There are also very few resources to guide the end-to-end audit process, with most practitioners making use of a custom curated toolkit of often independently developed resources (raji_closing_2020, 80).
However, our participants’ aspirations for tooling suggest another path may be possible—a path towards lasting infrastructure that gives auditors additional levers to hold model operators accountable. In the remainder of this section, we suggest several ways that researchers, practitioners, policymakers and other stakeholders can bridge the gap between available tooling and practice, including the development of educational resources and tool selection frameworks, greater commitments to audit tool maintenance, centralized resource repositories, and funding commitments.
5.2.1. Implications for HCI
  • Shared educational resources. HCI practitioners are particularly well equipped to develop the education-first practices and training resources that may be helpful to facilitate a shift towards more informed and nuanced decision-making. Because stakeholders involved in auditing come from various backgrounds, tools and practices are likely to require some degree of education and training for those who come from non-engineering backgrounds in order to use and embrace emerging best practices. Also, education-first tooling can help lower the barrier for participation for affected communities.
  • Decision-support frameworks for AI audit tool selection. Auditors sought more than just a plethora of auditing tools; they also wished for robust frameworks to aid practitioners in choosing the right tool for the specific problem they encounter. Tool selection was a major source of uncertainty for practitioners; as one expert suggested, “it’s not just about creating a multitude of auditing tools but also about fostering decision support frameworks that empower practitioners to make informed choices based on the context they are dealing with” (P25). Future research could explore frameworks that assist audit practitioners in identifying and choosing between tools at each stage in an audit.
@@ 5.2.1. Implications for HCI
  • Institutionalized AI audit tool maintenance. Currently, the burden of building and maintaining tools to address technical debt falls on the auditors, raising concerns about the sustainability of auditing efforts. This maintenance challenge leads to questions about responsibility—who should develop and sustain the necessary tools and infrastructure to support auditing activities effectively? One participant suggested putting expiration dates on tools to avoid future inaccuracies (P7). Researchers and practitioners should ensure that newly developed tools are accompanied with long-term maintenance plans, possibly through collaborations with civil society organizations such as the Linux Foundation, which hosts several of the open source tools we found (IBM’s AI Fairness 360 (bellamy_ai_2018, 9), IBM’s AI Explainability 360, and the Adversarial Robustness Toolbox).
5.2.2. Broader Implications
  • Centralized AI audit resources. Participants unanimously agreed on the need for shared infrastructure that supports the auditing process. One participant commented that tools are “a superpower for journalists, and something that really is the future of accountability… There really needs to be some sort of public infrastructure [for auditing online platforms]” (P7). As it stands, the audit tooling landscape is widely decentralized, which makes it difficult for auditors to find and use multiple audit tools to develop more end-to-end methodologies. A centralized repository—accompanying regulatory guidance, for example—could house a set of various tools, guidelines, and procedures tailored to the auditing process. Such a repository would provide multiple benefits: easier tool discovery and selection, ultimately leading to increased awareness, accessibility, and language sharing; enhanced collaboration between AI audit tool practitioners and tool developers by informing auditors of existing tools, which may deter unnecessary tool duplication; and coordination between audit standards and audit tool standards, which could help mitigate concerns about audit integrity.
  • Funding. Multiple participants noted the value of long-term funding to support tool development and maintenance: “We need more funding for this space… especially when it comes to infrastructure” (P20). One participant hoped to find “funders who are on board for longer tool maintenance projects” (P7). Another suggested that funders may be incentivized to address infrastructure gaps by connecting “performance issues and ML to downstream business KPIs,” (P4) but noted the difficulty of doing so when talking about “nebulous things, like fairness and bias” (P4). Developing high quality audit tools—even when adapting existing open source tools—took resources. Authors of audit legislation—as well as other policymakers and foundations—should set aside funding and resources for these kinds of long-term projects to ensure that audit tools are high quality, long-lasting, and address the wide range of auditor needs identified in this study.
@@ 6. Conclusion
Ideally, AI audit studies will translate into tangible outcomes of accountability, but this outcome is far from a given reality. In order for the audit process to truly be feasible and effective, we—researchers, policymakers, regulators, and audit practitioners—need to invest in the infrastructure required for accountability. This will require a full effort on multiple fronts, including everything from the design and development of new tools; to new community infrastructure, communication standard-setting; to considering advocacy for certain policy positions. In the end, we cannot just accept the minimum expectation when it comes to auditing, and must push the boundaries of this work until it becomes the meaningful mechanism of accountability it promises to be.
@@ Acknowledgements.
Thanks to our participants for volunteering their time and insight. Thanks also to seminar participants at Carnegie Mellon University for their feedback on an earlier version of this work.
Appendix A Methods
A.1. Initial Sources
We drew our initial list of tools from:
@@ A.1. Initial Sources
  • tools mentioned in previous surveys of fairness and other toolkits (deng_exploring_2022, 30, 58, 46)
  • academic papers presenting new tools surfaced in a recent literature review of audit studies (birhane_ai_2024, 14), including
    • the last five years of conference proceedings from: FAccT, AIES, EAAMO, AAAI, CVPR, ICWSM, WWW, WACV, EECV, and the ACL Anthology; also, the ACM Digital Library (including CHI and IC2S2) with the terms “audit”, “accountability”, “case study”, “bias”, “fairness”, or “assurance” in the title or abstract;
    • reports from the government agencies ICO and NIST;
  • a convenience seeding (accounting for 79 of the initial 148 tools) of other prominent audit tooling projects, academic papers, and tool-building organizations that we had encountered in our work as researchers and practitioners.
A.2. Theoretical Sampling
Category descriptors from the initial taxonomy that were sourced either from our labels or from descriptors used by tools already collected were then searched in combination with general keywords we identified that were commonly used in the algorithmic auditing domain. These keywords include the following: “tool”, “audit”, “algorithm”, “accountability”, “responsible AI”, “fairness”, “discrimination”, and “AI”. For each of our categories, we conducted English Google searches using each keyword combined with each of our initial category descriptors (e.g., “participatory audit”, “participatory tool”, etc.). In order to determine a point of saturation for each search, we examined each Google search page of 30 results at a time until two pages in a row contained no references to specific audit tools. It is important to note that we conducted searches for all categories in our initial taxonomy, but focused on less saturated categories (for example, harms discovery and tools using participatory methods) in order to provide well-rounded definitions of categories that were less common and/or visible. Categories with a more saturated set of examples were not given an equal amount of additional sourcing.
In addition to these targeted Google searches, we also added several new sources of tools to our initial list:
@@ A.2. Theoretical Sampling
  • the Participatory ML Workshop at ICML (kulynych_participatory_2020, 54);
@@
Appendix B Interview Protocol
@@
Appendix C Additional Figures
(a) Distribution of forks, for tools with Github repository.
@@ Appendix C Additional Figures
Figure C.1. Usage statistics by taxonomy category. Sorted by type of organization (our classification). Box-and-whisker plots included for categories with more than 10 points.
Figure C.2. Number of tools by taxonomy category, sorted by type of organization (our classification).
@@
Figure C.2. Number of tools by taxonomy category, sorted by type of organization (our classification).
Figure C.3. Number of tools with code in each taxonomy stage.
@@
Figure C.3. Number of tools with code in each taxonomy stage.
Figure C.4. Number of tools by taxonomy category sorted by license type.
@@
Figure C.4. Number of tools by taxonomy category sorted by license type.
Figure C.5. Number of tools by taxonomy category sorted by audit target.
@@
Figure C.5. Number of tools by taxonomy category sorted by audit target.
Figure C.6. Number of tools by taxonomy category sorted by format.
@@
Figure C.6. Number of tools by taxonomy category sorted by format.
Figure C.7. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Sorted by type of organization (our classification). Tools from organizations without Crunchbase entries excluded.
@@
Figure C.7. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Sorted by type of organization (our classification). Tools from organizations without Crunchbase entries excluded.
Figure C.8. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Tools from organizations without Crunchbase entries excluded.
In this bar graph, we present the number of tools categorized by taxonomy. Additionally, we’ve organized the data based on the type of license and number of employees associated with each tool. The graph visually represents the quantity of tools within each taxonomy category, with bars that are further divided or stratified based on their respective licenses and number of employee.
Figure C.8. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Tools from organizations without Crunchbase entries excluded.
@@
Figure C.9. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Tools from organizations without Crunchbase revenue estimates excluded.
@@
In this bar graph, we present the number of tools categorized by taxonomy. Additionally, we’ve organized the data based on the type of organization and estimated revenue associated with each tool. The graph visually represents the quantity of tools within each taxonomy category, with bars that are further divided or stratified based on their respective type of organization and estimated revenue
Figure C.9. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Tools from organizations without Crunchbase revenue estimates excluded.
Figure C.10. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Tools from organizations without Crunchbase revenue estimates excluded.
In this bar graph, we present the number of tools categorized by taxonomy. Additionally, we’ve organized the data based on the type of license and estimated revenue associated with each tool. The graph visually represents the quantity of tools within each taxonomy category, with bars that are further divided or stratified based on their respective type of license and estimated revenue
Figure C.10. Number of tools by taxonomy category. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_nodate, 25). Tools from organizations without Crunchbase revenue estimates excluded.
@@
Figure C.11. Private (pre-IPO) organizations with Crunchbase entries ($114/311$ organizations). Total funding sourced from Crunchbase (noauthor_crunchbase_nodate, 25).
@@
Boxplot showing distribution of tools by organization type and total funding in millions USD
Figure C.11. Private (pre-IPO) organizations with Crunchbase entries ($114/311$ organizations). Total funding sourced from Crunchbase (noauthor_crunchbase_nodate, 25).
@@
Table 3. Top-funded audit tool builders, per Crunchbase data. Includes only private (pre-IPO) organizations with Crunchbase entries.
@@
Organization Total funding (millions USD) Estimated revenue Employees
Stages
OpenAI 11303.12 \$50M to \$100M 501-1000
Perf. Analysis, Trans. Infra., Standards
Databricks 3497 \$500M to \$1B 5001-10000
Trans. Infra., Perf. Analysis
OneTrust 1120 \$100M to \$500M 1001-5000
Harms Disc.
DataRobot 1000.598 \$100M to \$500M 501-1000
Perf. Analysis
Hugging Face 395.2 101-250
Perf. Analysis, Standards
GitHub 350 \$100M to \$500M 1001-5000
Perf. Analysis
H2O.ai 251.099999 \$10M to \$50M 251-500
Perf. Analysis
Weights & Biases 250 \$10M to \$50M 251-500
Perf. Analysis
HackerOne 159.4 \$10M to \$50M 1001-5000
Harms Disc.
Bugcrowd 78.65 \$1M to \$10M 501-1000
Harms Disc.
Arthur 60.3 51-100
Standards, Perf. Analysis
Pymetrics 56.63 \$1M to \$10M 51-100
Perf. Analysis
Fiddler AI 45.2 \$1M to \$10M 11-50
Perf. Analysis
TruEra 42.284998 51-100
Perf. Analysis
CognitiveScale 40 \$10M to \$50M 51-100
Perf. Analysis
Calypso AI 38.2 \$500M to \$1B 11-50
Standards, Perf. Analysis
Seldon 33.691771 \$1M to \$10M 51-100
Perf. Analysis
Open Data Institute 32.835579 \$1M to \$10M 11-50
Standards
LexisNexis 30 \$1B to \$10B 10001+
Advocacy
The Markup 20 \$1M to \$10M 11-50
Data Coll.
@@
Table 4. 20 most popular Github repositories for tools in our database, sorted by number of forks.
Organization Forks Issues Stars Stages
Scrapy 10140 714 48351 Data Coll.
Selenium 7762 222 27630 Data Coll.
Appium 5916 211 16867 Data Coll.
CARLA 3013 858 9592 Data Coll.
SHAP 2993 1463 20021 Perf. Analysis
Evals 2242 107 11684 Perf. Analysis
LIME 1762 97 10813 Perf. Analysis
Adversarial Robustness Toolbox 1047 128 3930 Perf. Analysis
Seldon Core 780 103 3882 Perf. Analysis
AI Fairness 360 715 165 2116 Perf. Analysis
Interpret 686 65 5674 Perf. Analysis
Big Bench 529 79 2223 Perf. Analysis
Tensorflow Privacy 433 107 1802 Perf. Analysis
Foolbox 418 37 2539 Perf. Analysis
CodeSearchNet 369 14 1992 Perf. Analysis
Fairlearn 360 179 1635 Perf. Analysis
Language Interpretability Tool 336 82 3189 Perf. Analysis
AI Explainability 360 305 55 1389 Perf. Analysis
Tensorflow Model Analyzer 265 33 1213 Perf. Analysis
Error Analysis 254 86 944 Perf. Analysis
Responsible AI Toolbox 254 86 944 Perf. Analysis

### v3

@@ 5.2. Moving beyond ad hoc toolkits, towards shared infrastructure
Our determination of the performance of the algorithm carries a lot of weight within the organization. It’s not like somebody else could just throw it and be like “Oh, we’ll go ask somebody else”… because we built the infrastructure, so we have that lever… If you can control the data sources or the ways to integrate algorithms into the data sources, that gives you power. (P3)
The external auditors we interviewed were especially skeptical of the data provided from these tools and aware of their unreliability, citing the shutdowns of transparency infrastructures like Reddit and Twitter’s APIs (P24) as well as Facebook’s CrowdTangle tool (P12). Open source tools (e.g., for data scraping) can sometimes fill the gap, but do not always cover the scope and complexity of practical audit work.
Our participants’ aspirations for tooling envision another path—a path towards lasting public infrastructure that gives auditors additional levers to hold model operators accountable (marda_public_2024). Directions for HCI research and development include:
  • Tool catalogs & other shared infrastructure. Tool selection was a major source of uncertainty for practitioners; as one expert suggested, “it’s not just about creating a multitude of auditing tools but also about fostering decision support frameworks that empower practitioners to make informed choices based on the context they are dealing with” (P25). In addition to HCI work evaluating the efficacy of audit tools (lee_landscape_2021; deng_exploring_2022; wong_seeing_2023; kaye_risky_2023; berman_scoping_2024), future research could explore frameworks that assist audit practitioners in identifying and choosing between tools at each stage in an audit. Policymakers can invest in these frameworks and publish catalogs of vetted tools for auditors to reference, similar to the OECD’s Tools for Trustworthy AI list (oecd_tools_2021). In general, shared AI inventories, AI incident databases, AI audit report registries, tool catalogs, regulatory guidance, and other centralized repositories or common transparency infrastructure could increase awareness, accessibility, and knowledge sharing, particularly if the audit community and affected stakeholders are empowered to not only utilize this infrastructure but also contribute to its development.
@@
  • Institutionalized tool maintenance & funding. Currently, the burden of building and maintaining tools to address technical debt falls on the auditors, raising concerns about the sustainability of auditing efforts. One participant suggested putting expiration dates on tools to avoid future inaccuracies (P7). They also hoped to find “funders who are on board for longer tool maintenance projects” (P7), but another noted that attempts to raise support by connecting “performance issues and ML to downstream business KPIs” (P4) was difficult when talking about “nebulous things, like fairness and bias” (P4). Policymakers and foundations should set aside funding and resources for long-term tooling projects to ensure that audit tools are high quality, long-lasting, and address the wide range of needs identified in this study. Some initiatives, such as the Mozilla Technology Fund (noauthor_mozilla_2024), the U.K. AI Safety Institute’s Systemic AI Safety Grants (noauthor_fast_2024), and the French AI Action Summit Public Interest fund (ai_action_summit_public_2024) are already positioned to make these investments. Researchers and practitioners should include long-term maintenance plans with newly developed tools, possibly in collaboration with civil society organizations such as the Linux Foundation, which hosts several of the open source tools we found (IBM’s AI Fairness 360 (bellamy_ai_2018), for example).
5.3. Ongoing Impact
This work was completed as part of the Open Source Audit Tooling (OAT) project at the Mozilla Foundation and has already had demonstrable impact in policy engagement and funding. After submitting public comments, these findings were cited several times in the U.S. National Telecommunication and Information Administration (NTIA) “Artificial Intelligence Accountability Policy Report” (goodman_ntia_2024), the “Summary report on the call for evidence on the Delegated Regulation on data access” for the E.U. Digital Services Act (leerssen_digital_2023), and the U.K. AI Safety Institute’s “International AI Safety Report” (bengio_international_2025). OAT team members presented these findings to regulators at the U.K.’s OfCom and the U.S. Federal Trade Commission. OAT team members also participated in advising the selection of two rounds of Mozilla Technology Fund (MTF) awardees, including 5 teams in an inaugural cohort (2022), and 8 teams in a follow-up cohort (2023) focused on AI audit tooling.
@@ 6. Conclusion
Ideally, AI audit studies will translate into tangible outcomes of accountability, but this outcome is far from certain. In order for the audit process to truly be feasible and effective, we—researchers, policymakers, and audit practitioners—need to invest in the infrastructure required for accountability. This will require a full effort on multiple fronts, including everything from the design and development of new tools; to new community infrastructure, communication standard-setting; to considering advocacy for certain policy positions. We cannot accept the minimum from AI auditing—we must push the boundaries of this practice until it becomes the meaningful mechanism of accountability it has the potential to be.
@@ Acknowledgements.
Thanks to our participants for volunteering their time and insight. Thanks also to seminar participants at Carnegie Mellon University and the Northeast HCI Meeting for their feedback on an earlier version of this work. This work was supported by the Mozilla Foundation. A.B. is supported by Science Foundation Ireland via the ADAPT Centre of Digital Content Technology funded under the European Regional Development Fund (ERDF) through Grant #13/RC/2106_P2. B.V. was supported by the University of Notre Dame IBM Tech-Ethics Lab.
[…]
Appendix A Reflections
We consider our own cultural and professional perspectives throughout the interviews and our analysis (charmaz_constructing_2014). In particular, we view our position as both external to prominent AI developers and deployments, but still situated within the Western AI industry as a critical consideration. Our project was financially supported by a prominent U.S.-based foundation, and all the authors are either graduate students or graduates of well-funded universities in the U.S. and Europe. We thus recognize that our analysis is primarily scoped to the U.S. and the E.U. and may not be representative of the global AI auditing landscape or appropriate for or informed by other contexts. Because we drew our initial sources from our own fieldwork as audit practitioners and used English search engines for theoretical sampling, our dataset consists primarily of English language tools from from Western organizations, and our taxonomy reflects our own particular position. [^10] We also attempted to run translated queries in non-English search engines (e.g., Baidu) and to add regional keywords (e.g., African) to our searches, but we were unable to find any additional tools with initial tests of these methods. It thus seems likely that non-English, non-Western tools exist that our theoretical sampling was unable to identify. Also, although we intentionally defined tools as resources more broadly and attempted to leave space for non-technical solutions in our analysis and discussion, we struggled to avoid a techno-solutionist framing in our conclusions (wong_seeing_2023). We hope that future work will expand and look outside of the mostly technical, Western perspective emphasized in this work.
Appendix B Glossary
Definition B.1.
Accountability: bovens_analysing_2007 defines accountability as “a relationship between an actor and a forum, in which the actor has an obligation to explain and to justify his or her conduct, the forum can pose questions and pass judgment, and the actor may face consequences.” We use the term “accountability” in this legal-political sense, meaning to face consequential judgment for system behaviors and impacts that do not align with articulated expectations and standards.
Definition B.2.
AI Audit: birhane_ai_2024 define an audit as “any independent assessment of an identified audit target via an evaluation of articulated expectations with the implicit or explicit objective of accountability”. Independence, as outlined by birhane_ai_2024, ensures the auditors (which may or may not belong to the same organization as the developers) are operationally distinct from the team that engineered the examined AI system, maintaining separation from the engineering process. Identified audit targets refer to concrete, specific objects of examination, ideally tied to real-world AI deployments or widely-used open-source algorithms or datasets, which serve as relevant proxies. Implicit or explicit accountability refers to audits designed to inform consequential judgments, measuring the deployment’s behaviors and impacts against clearly articulated expectations.
Definition B.3.
AI Auditor: An entity executing an AI audit, which may be viewed as either internal or external to the organization developing and/or operating the audited system. (Note that auditor independence is a nuanced spectrum, and there is not always a sharp divide between internal and external auditing.
  • Internal Auditor: raji_outsider_2022 define an internal auditor as an entity executing an audit or investigation with some contractual relationship with the audit target. They typically seek to minimize corporate liability and test for compliance with corporate or industry-wide expectations. Internal auditors are often hired voluntarily or to meet regulatory mandates.
  • External Auditor: raji_outsider_2022 define an external auditor as an entity executing an audit or investigation without any contractual relationship with the audit target. They typically execute audits voluntarily with a broader mandate of identifying and minimizing the harm impacting their constituents.
Definition B.4.
AI Audit Tool: We use the term AI audit tool to refer broadly to software, interfaces, code, benchmarks, frameworks, and other artifacts used by auditors in the AI audit process. Audit tools include resources that support algorithmic analysis and inspection (e.g., benchmarks/datasets, documentation templates) as well as resources that support the assessment of internal and external expectations for institutions across stages of design and development.
Definition B.5.
Abandonment & disgorgement: In the context of algorithmic systems and regulatory enforcement, the terms algorithm abandonment (johnson_fall_2024) and algorithm disgorgement (li_algorithmic_2022) refer to the destruction or abandonment of a system deemed harmful, unethical, or in violation of societal or regulatory standards. Abandonment refers to “an organization’s decision to stop designing, developing, or using an algorithmic system due to its (potential) harms” (johnson_fall_2024), while disgorgement further requires the deletion of both improperly obtained data and any machine learning models, algorithms, or outputs derived from such data (li_algorithmic_2022).
Appendix C Additional Methods
C.1. Initial Sources
We drew our initial list of 143 tools from:
@@ C.1. Initial Sources
  • tools mentioned in previous surveys of fairness and other toolkits (deng_exploring_2022; lee_landscape_2021; hickock_ethical_2023)
  • academic papers presenting new tools surfaced in a recent literature review of audit studies (birhane_ai_2024); this literature review included audit studies from
    • the last five years of conference proceedings from: FAccT, AIES, EAAMO, AAAI, CVPR, ICWSM, WWW, WACV, EECV, and the ACL Anthology; also, the ACM Digital Library (including CHI and IC2S2) with the terms “audit”, “accountability”, “case study”, “bias”, “fairness”, or “assurance” in the title, terms commonly used in AI studies published in computing venues;
    • reports from the government agencies ICO and NIST
  • a convenience sample (accounting for 79 of the initial 143 tools) of other prominent audit tooling projects, academic papers, and tool-building organizations that we had encountered in our work as researchers and audit practitioners; most of these tools also appear in the sources above.
We included both tools designed specifically for AI auditing (such as AI fairness toolkits) and generic tools that have been used in AI audits (such as Selenium (noauthor_selenium_2023)). Our dataset is not an exhaustive list of all tools that have been or could be used in AI audit practice. Note that there are substitutes and competitors for many of the tools in our dataset (such as qualitative coding software), but we did not include them unless we observed them in one of the sources above or our subsequent searches (§ C.2). Conversely, inclusion of a tool does not indicate the authors’ endorsement or preference.
C.2. Theoretical Sampling
Category descriptors from the initial taxonomy that were sourced either from our labels or from descriptors used by tools already collected were then searched in combination with general keywords we identified that were commonly used in the algorithmic auditing domain. These keywords include the following: “tool”, “audit”, “algorithm”, “accountability”, “responsible AI”, “fairness”, “discrimination”, and “AI”. For each of our categories, we conducted English Google searches using a selection of one or more relevant keywords combined with each of our initial category descriptors (e.g., “participatory audit”, “participatory tool”, etc.). In order to determine a point of saturation for each search, we examined each Google search page of 30 results at a time until two pages in a row contained no references to specific audit tools. It is important to note that we conducted searches for all categories in our initial taxonomy, but focused first on less saturated categories (for example, harms discovery and tools using participatory methods) in order to provide well-rounded definitions of categories that were less common and/or visible. Categories with a more saturated set of examples were not given an equal amount of additional sourcing.
In addition to these targeted Google searches, we also added several new sources of tools to our initial list based on our initial taxonomy:
@@ C.2. Theoretical Sampling
  • the Participatory ML Workshop at ICML (kulynych_participatory_2020);
@@
Appendix D Interview Protocol
@@
Appendix E Additional Landscape Analysis
To analyze the qualities of tools across our taxonomy, we manually labeled each tool with several tags describing the tool’s documentation and function: license (open-source or proprietary); organization type (for-profit, non-profit, government, or academic); intended audit target (automated decision system, online platforms, large pre-trained online platforms autonomous vehicles, and/or other); intended user (internal and/or external); and format (e.g. API, software product, code/data repository, white paper, and/or other). One author created the labels and at least one other author reviewed each label for agreement.
We also supplemented our dataset with data from Crunchbase (noauthor_crunchbase_2023) accessed in September 2023, a platform for tracking funding, employment, revenue and other data for technology ventures, and Github, a platform for hosting and developing software. From Github, we scraped repository activity—primarily the number of forks, stars, and issues—for the 98 tools with Github repositories in our dataset. Of the 347 organizations in our dataset, we were able to access Crunchbase records for 202 (accounting for 270 tools); 173 of these records include estimated employee counts. Of the 132 entries that are not for universities or government agencies, 91 include revenue estimates. We also collected total venture funding (adjusted to U.S. dollars) for 48 firms out of the 112 firms that are still private (i.e., have not undertaken an initial public offering). Additionally, we used the Google Scholar API to annotate each academic reference of an identified tool with the most recent available citation count in September 2023.
Figure E.1. Number of tools by taxonomy category, sorted by type of organization (our classification). Tools may be used in multiple stages.
@@ Appendix E Additional Landscape Analysis
Figure E.2. Number of tools with code in each taxonomy stage. Tools may be used in multiple stages.
[…]
@@
Figure E.3. Number of tools by taxonomy category sorted by license type. Tools may be used in multiple stages.
[…]
@@
Figure E.4. Number of tools by taxonomy category sorted by audit target. Tools may be used in multiple stages.
[…]
@@
Figure E.5. Number of tools by taxonomy category sorted by format. Tools may be used in multiple stages.
[…]
@@
Figure E.6. Number of tools by taxonomy category. Tools may be used in multiple stages. Workforce size of creating organization sourced from Crunchbase (noauthor_crunchbase_2023). Sorted by type of organization (our classification). Tools from organizations without Crunchbase entries excluded.
[…]
@@
Figure E.7. Number of tools by taxonomy category. Tools may be used in multiple stages. Estimated revenue of creating organization sourced from Crunchbase (noauthor_crunchbase_2023). Tools from organizations without Crunchbase revenue estimates excluded.
[…]
@@
In this bar graph, we present the number of tools categorized by taxonomy. Additionally, we’ve organized the data based on the type of organization and estimated revenue associated with each tool. The graph visually represents the quantity of tools within each taxonomy category, with bars that are further divided or stratified based on their respective type of organization and estimated revenue.
@@
Figure E.8. Private (pre-IPO) organizations with Crunchbase entries ($114/311$ organizations). Total funding sourced from Crunchbase (noauthor_crunchbase_2023). Tools may be used in multiple stages.
[…]
@@
Boxplot showing distribution of tools by organization type and total funding in millions USD.
@@
Table E.1. Top-funded audit tool builders, per Crunchbase data (noauthor_crunchbase_2023). Includes only private (pre-IPO) organizations with Crunchbase entries.
@@
Organization Total funding (millions USD) Estimated revenue Employees Stages
OpenAI 11303.12 \$50M to \$100M 501-1000 Perf. Analysis, Transp. Infra., Standards
Databricks 3497 \$500M to \$1B 5001-10000 Transp. Infra., Perf. Analysis
OneTrust 1120 \$100M to \$500M 1001-5000 Harms Disc.
DataRobot 1000.598 \$100M to \$500M 501-1000 Perf. Analysis
Hugging Face 395.2 101-250 Perf. Analysis, Standards
GitHub 350 \$100M to \$500M 1001-5000 Perf. Analysis
H2O.ai 251.099999 \$10M to \$50M 251-500 Perf. Analysis
Weights & Biases 250 \$10M to \$50M 251-500 Perf. Analysis
HackerOne 159.4 \$10M to \$50M 1001-5000 Harms Disc.
Bugcrowd 78.65 \$1M to \$10M 501-1000 Harms Disc.
Arthur 60.3 51-100 Standards, Perf. Analysis
Pymetrics 56.63 \$1M to \$10M 51-100 Perf. Analysis
Fiddler AI 45.2 \$1M to \$10M 11-50 Perf. Analysis
TruEra 42.284998 51-100 Perf. Analysis
CognitiveScale 40 \$10M to \$50M 51-100 Perf. Analysis
Calypso AI 38.2 \$500M to \$1B 11-50 Standards, Perf. Analysis
Seldon 33.691771 \$1M to \$10M 51-100 Perf. Analysis
Open Data Institute 32.835579 \$1M to \$10M 11-50 Standards
LexisNexis 30 \$1B to \$10B 10001+ Advocacy
The Markup 20 \$1M to \$10M 11-50 Data Coll.
@@
Table E.2. 20 most popular Github repositories for tools in our database, sorted by number of forks.
Tool Forks Issues Stars Stages
Scrapy 10458 667 52280 Data Coll.
Selenium 8101 242 30129 Data Coll.
Appium 6052 137 18618 Data Coll.
CARLA 3562 1079 11075 Data Coll.
SHAP 3245 806 22426 Perf. Analysis
Evals 2568 130 14631 Perf. Analysis
LIME 1795 120 11492 Perf. Analysis
Language Model Evaluation Harness 1673 338 6312 Perf. Analysis
Adversarial Robustness Toolbox 1146 150 4738 Perf. Analysis
AI Fairness 360 827 199 2401 Perf. Analysis
Seldon Core 827 203 4334 Perf. Analysis
Interpret 726 105 6198 Perf. Analysis
Big Bench 582 107 2802 Perf. Analysis
Tensorflow Privacy 447 121 1915 Perf. Analysis
Foolbox 421 27 2713 Perf. Analysis
Fairlearn 410 163 1888 Perf. Analysis
Purple Llama 404 1 2476 Perf. Analysis
CodeSearchNet 384 14 2172 Perf. Analysis
Language Interpretability Tool 351 113 3453 Perf. Analysis
Error Analysis 344 86 1325 Perf. Analysis
Responsible AI Toolbox 344 86 1325 Perf. Analysis
Figure E.9. Github forks by taxonomy category (for tools with Github repositories), sorted by type of organization (our classification). Tools may be used in multiple stages. Box-and-whisker plots included for categories with more than 10 points.
[…]
In a boxplot graph, we present the distribution of Github forks by taxonomy category. We stratify each boxplot by the organization type of its creator and label outliers.
Figure E.10. Citations by taxonomy category (for tools with papers), sorted by type of organization (our classification). Tools may be used in multiple stages. Box-and-whisker plots included for categories with more than 10 points.
[…]
In a boxplot graph, we present the distribution of citations by taxonomy category. We stratify each boxplot by the organization type of its creator and label outliers.